Prev Next

SW&IT Meeting Minutes 2019-12-09

Attendees: Andrew, Carl, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191104
LR20191104 - Bec, Matthieu 80.0
LR20191104 - Holmberg, Carl 80.0
LR20191104 - Colson, Andrew 0.0
LR20191104 - Edmeades, David 48.0
LR20191104 - Sola, Igor 63.0
LR20191104 - Kubanek, Petr 100.7
LR20191104 - Hooper, Stephen 80.0
LR20191104 - Zhang, Yang 70.0
LR20191104 - Zhang, Xianyu 72.0
20191118
LR20191118 - Bec, Matthieu 64.0
LR20191118 - Holmberg, Carl 64.0
LR20191118 - Colson, Andrew 0.0
LR20191118 - Edmeades, David 64.0
LR20191118 - Sola, Igor 64.0
LR20191118 - Kubanek, Petr 24.0
LR20191118 - Hooper, Stephen 64.0
LR20191118 - Zhang, Yang 61.0
LR20191118 - Zhang, Xianyu 64.0
20191202
LR20191202 - Bec, Matthieu 40.0
LR20191202 - Holmberg, Carl 40.0
LR20191202 - Colson, Andrew 0.0
LR20191202 - Edmeades, David 40.0
LR20191202 - Sola, Igor 0.0
LR20191202 - Kubanek, Petr 38.0
LR20191202 - Hooper, Stephen 64.0
LR20191202 - Zhang, Yang 40.0
LR20191202 - Zhang, Xianyu 0.0

News from management

  • AI capture, new WR entry format

Previous Action Items

Owner Action Due Date Completion Date
MB Present AI new scheme 2019-12-09 2019-12-09

Open topics

  • AHS on hydrogen
  • kitchen area: ideas to propose?
  • github tip: watching repos

Team reports

PK

  • helped with AGW 4 testing
  • IT6940 - for C/C++ group, please read the IT description, it contains details for what to watch for when using vectors in multithread application
  • IT3229 - discussions with JPH
  • learned you can change base branch in a pull request (changed from master to 2019B in https://github.com/LBTO/tcs/pull/120)
  • PMC issue https://github.com/LBTO/tcs/pull/121
  • merged & closed rotator displays pull request
  • newGCS
    • gcs/bin/cliazcam can read configuration files for CCDs, connect to simulator or mountain depending on TCSLOCALE
    • #, * prefixes to address camera completed
    • delete, clear commands
  • IT3852 - looks similar to IT7380, code for better debugging of those problems
  • IT7695 - problem on client side, closed
  • IT7950 - AGW4 problem, ongoing
  • Ice/IIF problems - "iif" must resolve for Ice to tcs2 to work

DE

  • lbcwindows-devel XP activation ended. Shut down for the near future, Yang is looking to see if its function is supported by Windows 10
  • Synology still stuck on volume remove; support response time is glacial
    • Finally responded with enough clues for us to find and kill the offending processes, after which I was able to remove the volume.
    • Leroy did the swap Friday morning
  • Tyler requested the ability to clear Vault checkouts as part of his cleanups. According to Ketiv/Autodesk this will require his user account to be granted some level of administrative privilege. Go/No-go?
  • Mountain 4G bridging: how to draw out requirements from the users?

MB

  • Last week
    • OC, WPC, Management
    • completed safety trainings
    • AO retreat
    • AHS test with CH
    • LBTI alarms email tests
    • Attended SharkVIS IIF over VPN discussions. Answered very categorically: no go.
    • sql zabbix closer inspection
    • DMS select-all in page view
  • This week
    • ObsWs presentation to sciops
    • DMS hashed urls, issue at startup
    • AO data follow up
    • DMS zabbix cleanup, one way or another

CH

  • Last week:
    • Asked DE about receiving/accessing email from lbti addressed to lbti-alerts@lbto.org. It appears we're stuck with using Gmail's API. I've prototyped a script that works with cmholm@gmail.com, but I'd need someone to generate a cert to work with the swg-noreply account.
    • Completed and tested a TMS laser status script. This will need a different approach, probably with a reading from a power meter, since it is failure to close the laser shutter that leads to light leaks.
    • Added TMS laser and HBS pump status to ALS configuration files.
    • Restarted work on a DMS telescope temperature visualization page.
    • Began draft of rationale to add HBS coolant temp alert to AHS.
    • Completed safety training
  • This week:
    • Complete AHS rationale draft.
    • Demo initial DMS telescope temp viz page.
    • Coordinate with Steve Ertel re INDI interaction LBTI-AHS.
    • Build ARGOS arbitrator.

YZ

  • Working on new 64-bit C++ version TCSSendWavefront (the older C version has some issues with current IIF/Ice).
  • Working on Raritan KVM client installation for MODS (testing Jar file).
  • Working with Olga to review the current LBCFPIA code (need to remove any hard-coded path values, and possibly merge multiple current locations into one).
  • Will need to update the LBC skyflat wiki program (per Olga's request).
  • Working on rebuilding the AO-Cravat database.
  • Talked with Robert Reynolds about updating the VCAN documents.

XZ

Last week:
  • SOUL lab support
  • AO retreat
  • Power loss reaction (re-program MOXA)
  • Power backplan dicussion
This week:
  • AO datasystem routine check in
  • Power backplane dicussion
  • Re-program and test MOXA

AC

  • AGW4 troubleshooting
  • OAC config file automated git commit
  • ARGOS lite status GUI - tested and behaves as expected
  • TODO: ARGOS satellite closure audio warning
  • Vacation Thursday and Friday

IS

Last week:
  • OT & PIT Went over the open ITs and queue issues.
  • PIT
    • foundation work resubmit option to the PIT so there are no or to limit duplicates
  • miniQ
    • cleaned up
This week
  • PIT
    • ongoing work on resubmission logic
  • OT
    • go over open issues and make an estimate of work
  • LUCI
    • Issue loading mask, wrong information in header files. Need to figure out why is that.
  • Safety training
  • Set up my new laptop

SH

  • ovirt -- memory on 420, IDSDM issues, install, container Linux
  • docker registry -- SSL certs, NFS volumes
  • parts incoming
  • observatory tools first cut
  • move of VMWare into new vCenter failed in Tucson.

next:

  • rm507-5, rm507-4 work
  • robs disk backing
  • ovirt hyperconvergence
  • virtualize web1.mountain, kill the cluster.
    • move "catalogs" onto hosted machine
  • new hardware installs
Topic revision: r19 - 09 Dec 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback